OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle.

OpenStudy (anonymous):

Perimeter of a rectangle (P) : 2L +2W L=4 inches more than its width L=4+W therefore P=2(4+W) + 2W P=34 now you'll have 34 = 2(4+W) + 2W solve for the width (W) will ya?

OpenStudy (anonymous):